# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v3

@@ Abstract
A class of $2\times 2$ random-matrix models is introduced for which the Brody distribution is the exact eigenvalue spacing distribution. The matrix elements consist of constrained sums of an exponential random variable raised to various powers that depend on the Brody parameter. The random matrices introduced here differ from those of the Gaussian Orthogonal Ensemble (GOE) in three important ways: the matrix elements are not independent and identically distributed (i.e., not IID) nor Gaussian-distributed, and the matrices are not necessarily real and/or symmetric. The first two features arise from dropping the classical independence assumption, and the third feature stems from dropping the quantum-mechanical conditions that are imposed in the construction of the GOE. In particular, the hermiticity condition, which in the present model, is a sufficient but not necessary condition for the eigenvalues to be real, is not imposed. Consequently, complex non-Hermitian $2\times 2$ random matrices with real or complex eigenvalues can also have spacing distributions that are intermediate between those of the Poisson and Wigner classes. Numerical examples are provided for different types of random matrices, including complex-symmetric matrices with real or complex-conjugate eigenvalues.
@@ I Introduction
Despite its successful and widespread use, the Brody distribution has no theoretical foundation in the context of eigenvalue statistics (i.e., it is not derived from any physical or mathematical theory). Equation (3) is essentially a surmise, but because it is a surmise based on the well-established phenomenon of level repulsion, it is often described as being a “phenomenological” [36] or “semi-empirical” [27] distribution (“heuristic” [21] is another common descriptor). While the Brody distribution falls within the scope of random matrix theory (RMT) [11], its association with RMT has until now been groundless since there are no preexisting random-matrix models whose eigenvalue spacing distributions are given by Eq. (3). In fact, the $2\times 2$ model of Nieminen and Muche [37] is (to the author’s knowledge) the only model that has been put forward as a candidate; unfortunately the spacing distribution for their model is not exactly the Brody distribution. Indeed, a long-standing criticism of the Brody distribution is that it does not derive from any known random-matrix model. The existence question [37] is resolved in the current paper, which presents a class of $2\times 2$ random-matrix models that has the Brody distribution as its exact eigenvalue spacing distribution. The question of uniqueness is also addressed.
It is important to state at the outset that the random matrices defined in this paper do not have the properties of traditional GOE or (more generally) Wigner matrices; in particular, the matrix elements are dependent and (in general) non-identically distributed. This is not a flaw since independent and identically-distributed (IID) matrix elements are model assumptions employed in classical RMT that can be dropped [38, 39, 40, 41] without necessarily altering the classical RMT results obtained under IID conditions [42, 43, 44, 45, 46, 47, 48, 49]. This “spectral universality” under different non-IID conditions in theory applies only to large matrices, but as will be later shown using explicit examples, the elements of $2\times 2$ random matrices need not be IID to uphold result (2). To the author’s knowledge, a general and rigorous proof of this specific fact does not exist in the literature. One other important feature of the random matrices considered here is that they are not always real and/or symmetric. GOE matrices owe their real and symmetric structure to assumptions and conditions that originate from quantum mechanics. In contrast (and for generality), the random-matrix model defined in this paper has no built-in quantum-mechanical assumptions or conditions.
@@
The contents of the paper are as follows. In Sec. II, a general class of $2\times 2$ correlated random-matrix models is defined. The eigenvalue spacing distribution for this class of models is then derived in Sec. III with the result being the Brody distribution. Three simple sub-classes of models with varying matrix-element structures are defined and examined in Sec. IV; the significance of each sub-class is discussed therein and illustrative examples are provided. Additional examples involving complex matrices and eigenvalues are given in Sec. V. Some generalizations and extensions of the model introduced in Sec. II are considered in Sec. VI. Numerical results for selected real and complex examples from Secs. IV and V are then presented in Sec. VII. Comments (Sec. VIII), the Conclusion (Sec. IX), and several Appendices (A-D) complete the paper.
@@ II A General Class of $2\times 2$ Correlated Multivariate Models
$\mathcal{M}_{n}=\left(\begin{array}[]{cc}\displaystyle g_{1}(X)+\sum_{m=1}^{n+1}k_{m}Y^{p_{(m-1)}\over{q+1}}&~{}~{}\left(\displaystyle\sum_{m=n+2}^{2n+2}k_{m}Y^{p_{[m-(n+2)]}\over{q+1}}\right)h(T)\\[17.07182pt] \displaystyle{1\over h(T)}\left(\sum_{m=2n+3}^{3n+3}k_{m}Y^{p_{[m-(2n+3)]}\over{q+1}}\right)&~{}~{}~{}\displaystyle g_{2}(X)+\sum_{m=3n+4}^{4n+4}k_{m}Y^{p_{[m-(3n+4)]}\over{q+1}}\end{array}\right),$ (6)
where $m$ and $n$ are positive integers, $\{T,X,Y\}$ are independent random variables whose distributions will be specified below, $\{g_{1}(X),g_{2}(X)\}$ are functions of $X$ whose difference is a fixed (real or complex) number for all supported values of $X$ (i.e., $g_{1}(x)-g_{2}(x)=\eta\in\mathbb{C}$, $\forall x\in\text{supp}(X)$), $h(T)$ is any function of $T$ having the property $h(T)\neq 0$ for all supported values of $T$, $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ are a constrained set of (real or complex) constants, and $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ ($p_{0}=0$) are a constrained set of unequal real constants that can (but need not) have parametric dependencies. [^2] For example, some or all of $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ could depend on the parameter $q$. The distributions of $\{T,X,Y\}$ are as follows. Let the random variable $Y\sim\text{Exp}(\sigma_{\text{e}})$, that is, let $Y$ have an exponential distribution with scale parameter $\sigma_{\text{e}}>0$: [^3] In this paper, the common notation $f_{V}(v)$ will be used to refer to the p.d.f. of an arbitrary random variable $V$ with lower-case $v$ then denoting the density variable corresponding to $V$.
@@
In the vernacular of RMT, $Y$ has a “Poisson distribution” (in the density variable $y$) when $\sigma_{\text{e}}=1$. The random variables $X$ and $T$ can have any arbitrary continuous distributions $f_{X}(x)$ and $f_{T}(t)$, respectively (including exponential distributions). Thus, apart from the functions $\{g_{1}(X),g_{2}(X),h(T)\}$, the matrix elements are constrained sums of power-transformed exponential random variables. The matrix elements are furthermore subject to the condition that the discriminant of the characteristic polynomial of matrix $\mathcal{M}_{n}$, denoted here by $\mathcal{D}(\mathcal{M}_{n})$, is proportional to $Y^{2\over{q+1}}$:
@@
where the “discriminant constant” $k$ is a positive real constant. [^4] The constant $k$, which will depend on the constants $\{k_{m}\}$ in model (6), could also be negative, in which case, the eigenvalues are not real; this case will be considered in Sec. V.2. As will become clear in Sec. III, condition (8) is a sufficient condition for the eigenvalue spacing distribution of $\mathcal{M}_{n}$ to be the Brody distribution. It is important to note that, in general, the matrix elements of the above-defined model are not independent nor identically distributed.
@@ III Derivation of the Eigenvalue Spacing Distribution
Note that Eq. (10) is valid only when the eigenvalues are real, which will be the case for almost all sub-classes of models considered in this paper (exceptions occur in Sec. V.2).
@@
where $k>0$ is the discriminant constant. Note that $S$ is independent of the functions $\{g_{1}(X),g_{2}(X),h(T)\}$. As alluded to in Sec. II, the simple functional form for $S$ given by Eq. (11) is due to condition (8).
@@
$W\equiv Y^{1\over{q+1}}\sim\text{Weibull}\left(\theta=\sigma_{\text{e}},\tau=q+1\right),$ (12)
@@
$f_{S}(s)$ $=$ $f_{\sqrt{k}W}(s)={1\over\sqrt{k}}f_{W}\left(w={s\over\sqrt{k}};\theta=\sigma_{\text{e}},\tau=q+1\right)$ (13)
@@
$\mu_{S}=\mu_{\sqrt{k}W}=\sqrt{k}\mu_{W}(\theta=\sigma_{\text{e}},\tau=q+1)=\sqrt{k}\sigma_{\text{e}}^{1\over{q+1}}\Gamma\left({{q+2}\over{q+1}}\right).\quad$ (15)
@@
IV Three Sub-Classes of Models
Many different sub-classes of models can be constructed from the general multivariate model defined in Sec. II. For future reference, the term “cropped” shall be used to refer to models having all of the following properties: (i) $g_{1}(X)=g_{2}(X)=0$, (ii) $h(T)=1$, and (iii) no lone constants (in the summations); models having only properties (i) and (ii) will be referred to as “pruned”; models having properties (ii) and (iii) and (instead of (i)) the property $g_{1}(X)=g_{2}(X)\neq 0$ will be referred to as “trimmed”. Low-order sub-classes of models (e.g., $n=2$ or $n=3$) offer a degree of parametric versatility (via the existence of one independent parameter) and are algebraically simple. Higher-order models (e.g., $n>4$) can be more parametrically versatile but they are also more algebraically complicated. For example, imposing condition (8) on a cropped trivariate model
@@ IV Three Sub-Classes of Models
Low-order models are sufficiently versatile and algebraically simple and will hence be the focus of this paper. In the following subsections, three simple sub-classes of models referred to below as sub-classes A, B, and C, will be studied. As notational conveniences, the constrained constants $\left\{{p_{1}},{p_{2}},{p_{3}},\ldots\right\}$ appearing in the general model (6) will be denoted by $\left\{a,b,c,\ldots\right\}$ and the constants $\{k_{m}\}$ appearing in model (6) will be replaced by a simpler set of constants $\{c_{j}\}$. In this way, the sub-classes can be written as stand-alone models without explicit reference to the general model (6). To illustrate this notational scheme, consider the sub-class of
$\mathcal{M}_{2}=\left(\begin{array}[]{cc}\displaystyle g_{1}(X)+\sum_{m=1}^{3}k_{m}Y^{p_{(m-1)}\over{q+1}}&~{}~{}\displaystyle\left(\sum_{m=4}^{6}k_{m}Y^{p_{(m-4)}\over{q+1}}\right)h(T)\\[14.22636pt] \displaystyle{1\over h(T)}\left(\sum_{m=7}^{9}k_{m}Y^{p_{(m-7)}\over{q+1}}\right)&~{}~{}\displaystyle g_{2}(X)+\sum_{m=10}^{12}k_{m}Y^{p_{(m-10)}\over{q+1}}\end{array}\right)$ (26)
where the constants $\{k_{2},k_{3},k_{6},k_{8},k_{11},k_{12}\}$ are all identically zero and $h(T)=1$ (referred to here as sub-class E of class $\mathcal{M}_{2}$):
@@
The following three subsections consider simple trimmed, cropped, and pruned sub-classes of models that derive from the general model (6). The understanding is that these models are among the most basic $2\times 2$ (correlated) models possessing Brody spacing distributions and that (if desired) the random functions $\{g_{1}(X),g_{2}(X),h(T)\}$ can be judiciously added to incorporate additional randomization (i.e., reduce correlation) and/or impose independence between particular matrix elements.
IV.1 Sub-Class A
Sub-Class A consists of asymmetric random matrices of the (trimmed) form:
@@ IV.1.2 Significance of Sub-Class A
IV.1.3 Generalization of a Special Case
@@
$\mathcal{M}^{\text{(g)}}_{A}(c_{2}=0)=\left(\begin{array}[]{cc}X+c_{1}Y^{1\over{q+1}}&0\\ V&X+c_{4}Y^{1\over{q+1}}\end{array}\right),~{}~{}$ (41c)
$\mathcal{M}^{\text{(g)}}_{A}(c_{3}=0)=\left(\begin{array}[]{cc}X+c_{1}Y^{1\over{q+1}}&V\\ 0&X+c_{4}Y^{1\over{q+1}}\end{array}\right),~{}~{}$ (41f)
@@
IV.1.4 A Related Class of Models

### v4

# Current Address: ]Kinlin School, Fanshawe College, London, Ontario, Canada

@@ Abstract
A class of $2\times 2$ random-matrix models is introduced for which the Brody distribution is the exact eigenvalue spacing distribution. The matrix elements consist of constrained finite sums of an exponential random variable raised to various powers that depend on the Brody parameter. The random matrices introduced here differ from those of the Gaussian Orthogonal Ensemble (GOE) in three important ways: the matrix elements are not independent and identically distributed (i.e., not IID) nor Gaussian-distributed, and the matrices are not necessarily real and/or symmetric. The first two features arise from dropping the classical independence assumption, and the third feature stems from dropping the quantum-mechanical conditions imposed in the construction of the GOE. In particular, the hermiticity condition, which in the present class of models, is a sufficient but not necessary condition for the eigenvalues to be real, is not imposed. Consequently, complex non-Hermitian $2\times 2$ random matrices with real or complex eigenvalues can also have spacing distributions that are intermediate between those of the Poisson and Wigner classes. Numerical examples are provided for different types of random matrices, including complex-symmetric matrices with real or complex-conjugate eigenvalues. Various generalizations and extensions are discussed including a simple modification that effectuates cross-over transitions between other classes of eigenvalue spacing statistics. The case of a cross-over transition between semi-Poisson and Ginibre spacing statistics is presented as a novel example.
@@ I Introduction
Despite its successful and widespread use, the Brody distribution has no theoretical foundation in the context of eigenvalue statistics (i.e., it is not derived from any physical or mathematical theory). Equation (3) is essentially a surmise, but because it is a surmise based on the well-established phenomenon of level repulsion, it is often described as being a “phenomenological” [36] or “semi-empirical” [27] distribution (“heuristic” [21] is another common descriptor). While the Brody distribution falls within the scope of random matrix theory (RMT) [11], its association with RMT has thus far been groundless given the conspicuous absence of random-matrix models having Brody spacing distributions. In fact, the $2\times 2$ model of Nieminen and Muche [37] is (to the author’s knowledge) the only model that has been put forward as a candidate; unfortunately the spacing distribution for their model is not exactly the Brody distribution. Indeed, a long-standing criticism of the Brody distribution is that it does not derive from any known random-matrix model. The existence question [37] is resolved in the current paper, which presents a class of $2\times 2$ random-matrix models that has the Brody distribution as its exact eigenvalue spacing distribution. The question of uniqueness is also addressed.
It is important to state at the outset that the random matrices defined in this paper do not have the properties of traditional GOE or (more generally) Wigner matrices; in particular, the matrix elements are dependent and (in general) non-identically distributed. This is not a flaw since independent and identically-distributed (IID) matrix elements are model assumptions employed in classical RMT that can be dropped [38, 39, 40, 41] without necessarily altering the classical RMT results obtained under IID conditions [42, 43, 44, 45, 46, 47, 48, 49]. This “spectral universality” under different non-IID conditions in theory applies only to large matrices, but as will be later shown using explicit examples, the elements of $2\times 2$ random matrices need not be IID to uphold result (2). To the author’s knowledge, a general and rigorous proof of this specific fact does not exist in the literature. One other important feature of the random matrices considered here is that they are generally non-symmetric (when the matrices are real) and non-Hermitian (when the matrices are complex). GOE matrices owe their real and symmetric structure to assumptions and conditions that originate from traditional (i.e., Hermitian) quantum mechanics. In contrast (and for generality), the random-matrix model defined in this paper has no built-in quantum-mechanical assumptions or conditions.
@@
The contents of the paper are as follows. In Section II, a general class of $2\times 2$ correlated random-matrix models is defined. The eigenvalue spacing distribution for this class of models is then derived in Section III with the result being the Brody distribution. Terminology and notation for sub-classes are set out in Section IV. Three simple non-symmetric sub-classes of models with varying matrix-element structures are then defined and examined in Section V; the significance of each sub-class is discussed therein and illustrative examples are provided. Some simple symmetric sub-classes of models are presented in Section VI. Additional examples involving complex matrices and eigenvalues are given in Section VII. Some generalizations and extensions of the model introduced in Section II are considered in Section VIII. Numerical results for selected real and complex examples from Sections V and VII are then presented in Section IX. Comments (Section X), the Conclusion (Section XI), and several Appendices (A-D) complete the paper.
@@ II A General Class of $2\times 2$ Correlated Multivariate Models
$\mathcal{M}_{n}=\mathcal{A}\left(\begin{array}[]{cc}\displaystyle g_{1}(X)+\sum_{m=1}^{n+1}k_{m}Y^{p_{(m-1)}\over{q+1}}&~~\left(\displaystyle\sum_{m=n+2}^{2n+2}k_{m}Y^{p_{[m-(n+2)]}\over{q+1}}\right)h(T)\\[17.07182pt] \displaystyle{1\over h(T)}\left(\sum_{m=2n+3}^{3n+3}k_{m}Y^{p_{[m-(2n+3)]}\over{q+1}}\right)&~~~\displaystyle g_{2}(X)+\sum_{m=3n+4}^{4n+4}k_{m}Y^{p_{[m-(3n+4)]}\over{q+1}}\end{array}\right),$ (6)
where $m$ and $n$ are positive integers, $\{T,X,Y\}$ are independent random variables whose distributions will be specified below, $\{g_{1}(X),g_{2}(X)\}$ are functions of $X$ whose difference is a fixed (real or complex) number for all supported values of $X$ (i.e., $g_{1}(x)-g_{2}(x)=\eta\in\mathbb{C}$, $\forall x\in\text{supp}(X)$), $h(T)$ is any function of $T$ having the property $h(T)\neq 0$ for all supported values of $T$, $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ are a constrained set of (real or complex) constants, and $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ ($p_{0}=0$) are a constrained set of unequal real constants that can (but need not) have parametric dependencies. The prefactor $\mathcal{A}$ (outside the matrix) is an overall multiplicative factor for each of the matrix elements that can (in general) depend on any of the prior-defined constants or variables. The distributions of $\{T,X,Y\}$ are as follows. Let $Y\sim\text{Exp}(\sigma_{\text{e}})$, that is, let $Y$ have an exponential distribution with scale parameter $\sigma_{\text{e}}>0$: [^2] In this paper, the common notation $f_{V}(v)$ will be used to refer to the p.d.f. of an arbitrary random variable $V$ with lower-case $v$ then denoting the density variable corresponding to $V$.
@@
In the vernacular of RMT, $Y$ has a “Poisson distribution” (in the density variable $y$) when $\sigma_{\text{e}}=1$. The random variables $X$ and $T$ can have any arbitrary continuous distributions $f_{X}(x)$ and $f_{T}(t)$, respectively (including exponential distributions). Thus, apart from the functions $\{g_{1}(X),g_{2}(X),h(T)\}$ (and a possibly non-constant prefactor $\mathcal{A}$), the matrix elements are constrained sums of power-transformed exponential random variables. As such, the matrix elements are generally not independent nor identically distributed. Lastly, the discriminant of the characteristic polynomial of $\mathcal{M}_{n}$, denoted by $\mathcal{D}(\mathcal{M}_{n})$, must satisfy the condition:
@@
where the “discriminant constant” $k$ is a positive real constant. [^3] The constant $k$, which will depend on the constants $\{k_{m}\}$ in model (6), could also be negative, in which case, the eigenvalues are not real; this case will be considered in Section VII.2. As will become clear in Section III, condition (8) is a sufficient condition for the eigenvalue spacing distribution of $\mathcal{M}_{n}$ to be the Brody distribution.
@@ III Derivation of the Eigenvalue Spacing Distribution
Note that Eq. (10) is valid only when the eigenvalues are real, which will be the case for almost all sub-classes of models considered in this paper (exceptions occur in Section VII.2).
@@
where $k>0$ is the discriminant constant. Note that $S$ is independent of the functions $\{g_{1}(X),g_{2}(X),h(T)\}$. As alluded to in Section II, the simple functional form for $S$ given by Eq. (11) is due to condition (8).
@@
$W\equiv Y^{1\over{q+1}}\sim\text{Weibull}\left(\kappa=\sigma_{\text{e}},\tau=q+1\right),$ (12)
@@
$f_{S}(s)$ $=$ $f_{\sqrt{k}W}(s)={1\over\sqrt{k}}f_{W}\left(w={s\over\sqrt{k}};\kappa=\sigma_{\text{e}},\tau=q+1\right)$ (13)
@@
$\mu_{S}=\mu_{\sqrt{k}W}=\sqrt{k}\mu_{W}(\kappa=\sigma_{\text{e}},\tau=q+1)=\sqrt{k}\sigma_{\text{e}}^{1\over{q+1}}\Gamma\left({{q+2}\over{q+1}}\right).\quad$ (15)
@@
IV Sub-Classes of $\mathcal{M}_{n}$: Terminology and Notation
Many different sub-classes of models can be constructed from the general multivariate model defined in Section II. For future reference, the term “cropped” shall be used to refer to models having all of the following properties: (i) $g_{1}(X)=g_{2}(X)=0$, (ii) $\mathcal{A}=h(T)=1$, and (iii) no lone constants (in the summations); models having only properties (i) and (ii) will be referred to as “pruned”; models having properties (ii) and (iii) and (instead of (i)) the property $g_{1}(X)=g_{2}(X)\neq 0$ will be referred to as “trimmed”. Low-order sub-classes of models (e.g., n=2,3n=2,~3, or $4$) offer a degree of parametric versatility (via the existence of one independent parameter) and are algebraically simple. Higher-order models (e.g., $n>4$) can be more parametrically versatile but they are also more algebraically complicated. For example, imposing condition (8) on a cropped trivariate model
@@ IV Sub-Classes of $\mathcal{M}_{n}$: Terminology and Notation
Low-order models are sufficiently versatile and algebraically simple and will hence be the focus of this paper. In the following sections, several different sub-classes of models will be studied. As a notational convenience, the constrained constants $\left\{{p_{1}},{p_{2}},{p_{3}},\ldots\right\}$ appearing in the general model (6) will be denoted by $\left\{a,b,c,\ldots\right\}$ and the constants $\{k_{m}\}$ appearing in model (6) will be replaced by a simpler set of constants $\{c_{j}\}$. In this way, the sub-classes can be written as stand-alone models without explicit reference to the general model (6). To illustrate this notational scheme, consider the sub-class of
$\mathcal{M}_{2}=\mathcal{A}\left(\begin{array}[]{cc}\displaystyle g_{1}(X)+\sum_{m=1}^{3}k_{m}Y^{p_{(m-1)}\over{q+1}}&~~\displaystyle\left(\sum_{m=4}^{6}k_{m}Y^{p_{(m-4)}\over{q+1}}\right)h(T)\\[14.22636pt] \displaystyle{1\over h(T)}\left(\sum_{m=7}^{9}k_{m}Y^{p_{(m-7)}\over{q+1}}\right)&~~\displaystyle g_{2}(X)+\sum_{m=10}^{12}k_{m}Y^{p_{(m-10)}\over{q+1}}\end{array}\right)$ (26)
where the constants $\{k_{2},k_{3},k_{6},k_{8},k_{11},k_{12}\}$ are all identically zero and $\mathcal{A}=h(T)=1$ (referred to here as sub-class E of class $\mathcal{M}_{2}$):
@@
V Simple Non-Symmetric Models
Simple trimmed, cropped, and pruned sub-classes of models are (under condition (8)) typically non-symmetric. The following three subsections consider (in turn) a low-order example of each aforementioned type. The understanding is that these non-symmetric models are among the most basic $2\times 2$ (correlated) models possessing Brody spacing distributions and that (if desired) the random functions $\{g_{1}(X),g_{2}(X),h(T)\}$ can be judiciously added to incorporate additional randomization (i.e., reduce correlation) and/or impose independence between particular matrix elements.
V.1 Sub-Class A
Sub-Class A consists of random matrices of the (trimmed) form:
@@ V.1.2 Significance of Sub-Class A
V.1.3 Generalization of a Special Case
@@
$\mathcal{M}^{\text{(g)}}_{A}(c_{2}=0)=\left(\begin{array}[]{cc}X+c_{1}Y^{1\over{q+1}}&0\\ V&X+c_{4}Y^{1\over{q+1}}\end{array}\right),$ (41c)
$\mathcal{M}^{\text{(g)}}_{A}(c_{3}=0)=\left(\begin{array}[]{cc}X+c_{1}Y^{1\over{q+1}}&V\\ 0&X+c_{4}Y^{1\over{q+1}}\end{array}\right),$ (41f)
@@
V.1.4 A Related Class of Models